OpenStudy (anonymous):

How do I find the solution to this? x2 + 2x = -2 It is a complex solution. Do I complete the square?

jimthompson5910 (jim_thompson5910):

either that or use the quadratic formula (after you get that -2 to the right side by adding 2 to both sides)

jimthompson5910 (jim_thompson5910):

x^2 + 2x = -2 x^2 + 2x + 2 = -2 + 2 x^2 + 2x + 2 = 0 now use the quadratic formula

jimthompson5910 (jim_thompson5910):

I prefer the quadratic formula since it's more simple and direct in my opinion.

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome